OpenStudy (anonymous):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work. plz someone >~<

OpenStudy (campbell_st):

steps 1. subtract ut from both sides of the equation 2. multiply both sides of the equation by 2 3. divide both sides of the equation by t^2

OpenStudy (anonymous):

how would it look like? >< so d-ut=0.5at2 ? then ? d-ut2=0.5at?

OpenStudy (campbell_st):

the 1st step is correct but put it in brackets before multiplying by 2....

OpenStudy (anonymous):

d-ut(2)=0.5at? then d-ut(2)/t=A?

OpenStudy (campbell_st):

ok... 1st step correct \[d - ut = 0.5 at^2\] use brackets and multiply by 2 \[2(d - ut) = at^2\]

OpenStudy (anonymous):

oh ;-; so 2d-2ut=at2??

OpenStudy (campbell_st):

yes... but keep it written in brackets with 2 outside... now divide by t^2

OpenStudy (anonymous):

2(d-ut)=at2 2(d-ut)/t^2=A? ;-;

OpenStudy (campbell_st):

thats correct, well done

OpenStudy (anonymous):

wow thank you so much! !! im so happy ^_^

OpenStudy (anonymous):

@hartnn can you explain what the answer Is I don't fully understand